# Canonical decompositions of 3-connected graphs

### v2

@@
Abstract. We offer a new structural basis for the theory of 3-connected graphs, providing a unique decomposition of every such graph into parts that are either quasi 4-connected, wheels, or thickened $K_{3,m}$’s. Our construction is explicit, canonical, and has the following applications: we obtain a new theorem characterising all Cayley graphs as either essentially 4-connected, cycles, or complete graphs on at most four vertices, and we provide an automatic proof of Tutte’s wheel theorem.
2020 Mathematics Subject Classification. 05C40, 05C75, 05C83, 05E18
Key words. 3-connected, decomposition, canonical, wheel, 4-connected, 3-separation, tri-separation
${}^{\clubsuit}$niversity of Birmingham, Birmingham, UK, funded by EPSRC, grant number EP/T016221/1
@@ Introduction
The obvious guess how the solution might extend to 3-separators of 3-connected graphs is this: every 3-connected graph decomposes into 4-connected graphs, wheels and $K_{3}$’s, by cutting precisely at its totally-nested 3-separators. This guess turns out to be wrong, as the following three examples demonstrate.
@@
A mixed-separation of a graph $G$ is a pair $(A,B)$ such that $A\cup B=V(G)$ and both $A\setminus B$ and $B\setminus A$ are nonempty. We refer to $A$ and $B$ as the sides of the mixed-separation. The separator of $(A,B)$ is the disjoint union of the vertex set $A\cap B$ and the edge set $E(A\setminus B,B\setminus A)$. If the separator of $(A,B)$ has size three, we call $(A,B)$ a mixed 3-separation.
@@
Two mixed-separations $(A,B)$ and $(C,D)$ of $G$ are nested if, after possibly switching the name $A$ with $B$ or the name $C$ with $D$, we have $A\subseteq C$ and $B\supseteq D$. A tri-separation of $G$ is totally-nested if it is nested with every tri-separation of $G$. A tri-separation $(A,B)$ of a 3-connected graph $G$ is trivial if $A$ and $B$ are the sides of a 3-edge-cut with a side of size one. The tri-separations that we will use to decompose $G$ are the totally-nested nontrivial tri-separations of $G$.
@@
Given a 3-connected graph $G$ and a set $N$ of pairwise nested tri-separations, we can say which parts we obtain by decomposing $G$ along $N$. Roughly speaking, these are maximal subgraphs of $G$ that lie on the same side of every tri-separation in $N$, with some edges added to represent external connectivity in $G$. We call the resulting minors of $G$ the torsos of $N$, as they generalise the well known torsos of tree-decompositions from the theory of graph minors. See Section 2.2 for details.
@@
Let $G$ be a 3-connected graph and let $N$ denote its set of totally-nested nontrivial tri-separations. Each torso $\tau$ of $N$ is a minor of $G$ and satisfies one of the following:
@@
We emphasise that the sets $N=N(G)$ obviously are canonical, meaning that they commute with graph-isomorphisms: $N(\varphi(G))=\varphi(N(G))$ for all $\varphi\colon G\to G^{\prime}$. Our proof of Theorem 1 offers additional structural insights which can be used to refine Theorem 1; see Theorem 2.2.1. All graphs in this paper are finite or infinite unless stated otherwise; in particular, Theorem 1 includes infinite 3-connected graphs. It is not clear to us how the sets $N(G)$ could determine tree-decompositions, but we have a natural explanation for this: there exists a notion more general than tree-decomposition which can be used to express the sets $N(G)$; see Section 3.1.
@@
We give the precise definition of ‘essentially 4-connected’ in Section 1.4; the main difference to ‘quasi 4-connected’ is that we allow 3-edge-cuts that have a side which is equal to a triangle. Applications strengthens [37, Theorem 3.4.2], a classical tool in geometric group theory from the textbook of Godsil and Royle, in a special case.
Another application comes in the form of an automatic proof of Tutte’s wheel theorem [62]. In the upcoming work [17], Theorem 1 will be used to construct an FPT algorithm for connectivity augmentation from 0 to 4, and the property of total nestedness is crucial for that; see Chapter 3.
@@
    Recently, coverings as known from Topology have been employed to systematically construct graph-decompositions, tree-decompositions where the tree may be any arbitrary graph, by applying classical theorems about tree-decompostions to the covering of a graph, and then folding the tree-decomposition to a graph-decomposition [26]. For the Tutte-decomposition, it is known how to achieve this directly without employing coverings [7]. For the covering approach, canonicty is paramount: it makes the entire construction work. For the direct approach, the description of the Tutte-decomposition via total-nestedness is key. The construction from [26] can be generalised so that it can be applied to the tri-separation decomposition; we would be excited to see this happen.
@@
Our work complements existing work on decomposing graphs along separators of arbitrary size and the corresponding theory of tangles [2, 6, 9, 10, 11, 12, 16, 22, 23, 24, 25, 28, 29, 30, 39, 40, 42, 56]. It would be most exciting to try to extend our work to separators of larger size, see Chapter 3 for an open question in this direction. A fair share of the work on 3-connected graphs studies which substructures are ‘(in-)essential’ to 3-connectivity (in the sense that their contraction or deletion preserves 3-connectivity); this includes the work of Ando, Enomoto and Saito [3] and of Kriesell [43, 44, 45, 46, 47, 48]. The structure of 3-separations in matroids is a well-studied topic; a fundamental result in this area is the decomposition result of Oxley, Semple and Whittle [52]. It would be most natural to extend our ideas to matroids and this problem is discussed in Chapter 3. Hopes for a generalisation to directed graphs are fuelled by recent work of Bowler, Gut, Hatzel, Kawarabayashi, Muzi and Reich [4]. Our work is related to recent work of Esperet, Giocanti and Legrand-Duchesne [31], who employ Grohe’s techniques for 3-connected graphs to prove a general decomposition result of infinite graphs without an infinite clique-minor, and our result might provide an alternative perspective.
@@
An overview of the proof of Theorem 1 is given in Section 0.1. The remainder of the paper consists of three chapters and an appendix. Each chapter will feature its own comprehensive overview. In the first chapter, we introduce and prove the (Angry Tri-Separation Theorem). (1.1.1); this classifies the 3-connected graphs which have no totally-nested nontrivial tri-separations, and it will be the key ingredient of the proof of Theorem 1 as it deals with the special case $N(G)=\emptyset$. Applications can already be derived from the (Angry Tri-Separation Theorem)., so the first chapter also includes the proof of Applications. In the second chapter, we prove Theorem 1. The third chapter provides an outlook, which includes a discussion of the relation to tree-decompositions. Finally, the appendix offers a proof of the 2-separator theorem, but phrased in the language of this paper and with a structural strengthening added, for the sake of convenience and completeness.
0.1. Overview of the proof
Let $G$ be a 3-connected graph, and let $N$ denote the set of totally-nested nontrivial tri-separations of $G$. Let $\tau$ be an arbitrary torso of $N$. It is routine to verify that $\tau$ is 3-connected or a $K_{3}$, and that $\tau$ is a minor of $G$. So it remains to show that $\tau$ either is quasi 4-connected, a wheel, a thickened $K_{3,m}$ or $G=K_{3,m}$ with $m\geqslant 0$.
@@ 0.1. Overview of the proof
For the proof of (3), suppose that $\tau$ is lightly interlaced by a tri-separation $(A,B)$. Then we first note that the separator $S$ of $(A,B)$ consists of three vertices, and that $G\setminus S$ has at least four components. The four components ensure that $S$ is ‘4-connected’ in $G$, as every two vertices in $S$ can be linked by four internally vertex-disjoint paths in $G$ through the four components. The 4-connectivity of $S$ can then be used to show that every component $C$ of $G\setminus S$ determines a totally-nested tri-separation of $G$, with $C$ on one side, whose separator consists of vertices of $S$ or edges that join $C$ to $S$. The nontrivial ones amongst the totally-nested tri-separations determined by the components of $G\setminus S$ are precisely the ones that bound the torso $\tau$. If $G[S]$ is edgeless and all components of $G\setminus S$ have size one, then $G=K_{3,m}$ where $m$ is equal to the number of components of $G\setminus S$. Otherwise, a brief analysis shows that $\tau$ is a thickened $K_{3,m}$, where $m$ is equal to the number of components $C$ of $G\setminus S$ such that $|C|=1$ and $G[S]$ is edgeless.
Figure 1. The graph $G$ (on the right) can be constructed from a cycle $O$ (on the left) by replacing its bold edges by 2-connected graphs and adding the vertex $v$ together with its incident edges. Here $S(A,B)$ and $S(C,D)$ denote the separators of $(A,B)$ and $(C,D)$, respectively. They intersect in the vertex $v$. The cycle $O$ is highlighted in green.
@@
Figure 2. The graph $G$ together with the bold edges of $O$. For each bold edge of $O$ we indicate in grey its corresponding 2-connected replacement graph as given in the construction of $G$ from $O$. In this figure we refer to them as bags. The vertex $a$ is an endvertex of the bold edge $e$. Here the cycle $O$ has the pattern btx with regard to $e$ and $a$: the first letter b indicates that the edge $f_{1}$ on $O$ incident with $a$ aside from $e$ is bold, the second letter t indicates that the edge $f_{2}$ after that on $O$ is not bold (timid), and the letter x indicates that the edge $f_{3}$ after that can be arbitrary, bold or not. In our example, it is bold. For this pattern, there are three internally vertex-disjoint paths from $a$ to $v$: the first path $P_{1}$ connects to $v$ from the bag at $e$, the second path $P_{2}$ connects to $v$ from the bag at $f_{1}$, and the final path $P_{3}$ uses a path disjoint from $P_{2}-a$ through the bag at $f_{1}$, then traverses $f_{2}$ and eventually connects to $v$ from the bag at $f_{3}$.
@@
Having shown that the pseudo-reductions at the bold edges of $O$ are totally-nested, we know that they bound a torso $\tau^{\prime}$; this is not necessarily a torso of $N$, but of the set of pseudo-reductions. Using our knowledge of the structure of $G-v$ provided by $O$, and 3-connectedness of $G$, it is straightforward to show that $\tau^{\prime}$ is a wheel. So all that remains to show is that $\tau^{\prime}$ is equal to $\tau$. For this, we have a uniqueness-lemma, by which it suffices to show that no totally-nested nontrivial tri-separation of $G$ interlaces $\tau^{\prime}$. So we assume for a contradiction that some totally-nested nontrivial tri-separation $(U,W)$ of $G$ interlaces $\tau^{\prime}$. Roughly speaking, $(U,W)$ induces a mixed 2-separation of $O$. This induced mixed 2-separation cuts $O$ into two intervals. We carefully select a vertex or non-bold edge from each interval, and then add $v$ to obtain the separator of a mixed 3-separation $(E,F)$ of $G$. Then $(E,F)$ crosses $(U,W)$ by standard arguments, and with a bit of extra work we turn $(E,F)$ into a tri-separation that still crosses $(U,W)$, yielding a contradiction. Hence $\tau^{\prime}$ is equal to $\tau$. As we have already shown that $\tau^{\prime}$ is a wheel, the overview of the proof of (2) is complete.
Chapter 1 An angry theorem for tri-separations
1.1. Overview of this chapter
@@
This chapter provides the key ingredient for the proof of Theorem 1, which we call the (Angry Tri-Separation Theorem).. Essentially, it states that Theorem 1 holds in the special case where $N(G)$, the set of all totally-nested non-trivial tri-separations of $G$, is empty; that is: if $G$ is not itself quasi 4-connected, and if every nontrivial tri-separation of $G$ is crossed, then $G$ must be either a wheel or a $K_{3,m}$ for some $m\geqslant 3$.
1.1.1. Statement of the (Angry Tri-Separation Theorem).
@@
Here we give all the necessary definitions to then state the (Angry Tri-Separation Theorem). (1.1.1). An (oriented) mixed-separation of a graph $G$ is an ordered pair $(A,B)$ such that $A\cup B=V(G)$ and both $A\setminus B$ and $B\setminus A$ are non-empty. We call $A$ and $B$ the sides of $(A,B)$. The separator of $(A,B)$ is the disjoint union of the vertex set $A\cap B$ and the edge set $E(A\setminus B,B\setminus A)$. We denote the separator of $(A,B)$ by $S(A,B)$. The order of $(A,B)$ is the size $|S(A,B)|$ of its separator. A mixed-separation of order $k$ for $k\in{\mathbb{N}}$ is called a mixed $k$-separation for short. The separator of a mixed $k$-separation is a mixed $k$-separator. A mixed-separation $(A,B)$ of $G$ with no edges in its separator is called a separation of $G$. Separations of order $k$ are called $k$-separations and their separators are called $k$-separators.
@@
Figure 1.1. Separators of nontrivial tri-separations of the 4-wheel.
@@
Every nontrivial tri-separation in the 4-wheel, as found in Section 1.1.1, is crossed by another nontrivial tri-separation; see Figure 1.1. By contrast, every 3-cut with both sides of size at least two in a 3-connected graph determines a totally-nested nontrivial tri-separation (we will see this in Section 1.3).
@@
Theorem 1.1.1 (Angry Tri-Separation Theorem).
@@
1.1.2. Organisation of this chapter
In Section 1.2, we collect useful properties of tri-separations. In Section 1.3, we introduce tools that allow us to systematically study how tri-separations can cross. In Section 1.4, we deduce Applications from the (Angry Tri-Separation Theorem).. In Section 1.5, we employ the tools from the previous section to find necessary conditions for when a tri-separation is totally nested. In Section 1.6, we recall the 2-separation-version of Theorem 1, as we will need it in the proof of the (Angry Tri-Separation Theorem).. In Section 1.7, we will see why the 2-separation-version of Theorem 1 is helpful for finding totally-nested nontrivial tri-separations. In Section 1.8 and Section 1.9, we put together the tools developed in the previous sections and we complete the proof of the (Angry Tri-Separation Theorem)., where the former deals with special cases and the latter solves the general case.
1.2. Properties of tri-separations

### v3

@@
Abstract
We offer a new structural basis for the theory of 3-connected graphs, providing a unique decomposition of every such graph into parts that are either quasi 4-connected, wheels, or thickened $K_{3,m}$’s. Our construction is explicit, canonical, and has the following applications: we obtain a new theorem characterising all finite Cayley graphs as either essentially 4-connected, cycles, or complete graphs on at most four vertices, and we provide an automatic proof of Tutte’s wheel theorem.
[…]
[classification=text]
@@ Introduction
The obvious guess how the solution might extend to 3-separators of 3-connected graphs is this: every 3-connected graph decomposes into 4-connected graphs, wheels [^2] Every $k$-connected graph $G$ can be made $(k+1)$-connected by adding one more vertex and joining it to all original vertices. From this perspective, wheels are the 3-connected analogues of cycles. and $K_{3}$’s, by cutting precisely at its totally-nested 3-separators. This guess turns out to be wrong, as the following three examples demonstrate.
@@
A mixed-separation of a graph $G$ is a pair $(A,B)$ such that $A\cup B=V(G)$ and both $A\setminus B$ and $B\setminus A$ are nonempty. We refer to $A$ and $B$ as the sides of the mixed-separation. The separator of $(A,B)$ is the disjoint union of the vertex set $A\cap B$ and the edge set $E(A\setminus B,B\setminus A)$. If the separator of $(A,B)$ has size three, we call $(A,B)$ a mixed 3-separation. Mixed-separations can be viewed as separations of the barycentric subdivision of $G$.
@@
Two mixed-separations $(A,B)$ and $(C,D)$ of $G$ are nested if, after possibly switching the name $A$ with $B$ or the name $C$ with $D$, we have $A\subseteq C$ and $B\supseteq D$. A tri-separation of $G$ is totally-nested if it is nested with every tri-separation of $G$. A tri-separation $(A,B)$ of a 3-connected graph $G$ is nontrivial if both $G[A]$ and $G[B]$ contain a cycle; otherwise it is trivial. So $(A,B)$ is trivial if and only if $A$ and $B$ are the sides of a 3-edge-cut with a side of size one (Section 3). The tri-separations that we will use to decompose $G$ are the totally-nested nontrivial tri-separations of $G$.
@@
Given a 3-connected graph $G$ and a set $N$ of pairwise nested tri-separations, we can say which parts we obtain by decomposing $G$ along $N$. Roughly speaking, these are maximal subgraphs of $G$ that lie on the same side of every tri-separation in $N$, with some edges added to represent external connectivity in $G$. We call the resulting minors of $G$ the torsos of $N$, as they generalise the well known torsos of tree-decompositions from the theory of graph minors. We consider two types of torsos in this paper: compressed torsos and expanded torsos. What we call torso here will be called compressed torso later. See Section 2 for details; in particular Theorem 2.1.
While the block-cutvertex decomposition and the Tutte-decomposition can be described in terms of tree-decompositions, this is no longer evident for the tri-separation decomposition as the adhesion-sets of tree-decompositions cannot contain edges by definition. To overcome this, we introduce a notion of mixed-tree-decomposition that allows both vertices and edges in the adhesion-sets. Mixed-tree-decompositions sit in between tree-decompositions (that only allow vertices in adhesion-sets) and Wollan’s tree-cut decompositions [68] (that only allow edges in adhesion-sets). See Section 2 for details. Every mixed-tree-decomposition of a graph $G$ can be viewed as a tree-decomposition of the barycentric subdivision of $G$.
@@
Let $G$ be a 3-connected graph. Let ${{\mathcal{T}}}(G)$ denote the mixed-tree-decomposition of $G$ that is uniquely determined by the set $N(G)$ of all totally-nested nontrivial tri-separations of $G$. Each torso $\tau$ of ${{\mathcal{T}}}(G)$ is a minor of $G$ and satisfies one of the following:
@@
We emphasise that the sets $N(G)$ and mixed-tree-decompositions ${{\mathcal{T}}}(G)$ are obviously canonical, meaning that they commute with graph-isomorphisms: $N(\varphi(G))=\varphi(N(G))$ and ${{\mathcal{T}}}(\varphi(G))=\varphi({{\mathcal{T}}}(G))$ for all graph-isomorphisms $\varphi\colon G\to G^{\prime}$. Our proof of Theorem 1 offers additional structural insights which can be used to refine Theorem 1; see Theorem 2.1. All graphs in this paper are finite or infinite unless stated otherwise; in particular, Theorem 1 includes infinite 3-connected graphs.
@@
We give the precise definition of ‘essentially 4-connected’ in Section 5; the main difference to ‘quasi 4-connected’ is that we allow 3-edge-cuts that have a side which is equal to a triangle. A classical tool in geometric group theory from the textbook of Godsil and Royle [38, Theorem 3.4.2] states that every finite vertex-transitive $d$-regular graph is $k$-connected for $k:=\lceil\frac{2}{3}(d+1)\rceil$. Applications strengthens this tool in a special case.
Another application comes in the form of an automatic proof of Tutte’s wheel theorem [65]. In the upcoming work [18], Theorem 1 will be used to construct an FPT algorithm for connectivity augmentation from 0 to 4, and the property of total nestedness is crucial for that; see Section 7.
@@
    Recently, coverings as known from Topology have been employed to systematically construct graph-decompositions, tree-decompositions where the tree may be any arbitrary graph, by applying classical theorems about tree-decompostions to the covering of a graph, and then folding the tree-decomposition to a graph-decomposition [27]. For the Tutte-decomposition, it is known how to achieve this directly without employing coverings [8]. For the covering approach, canonicity is paramount: it makes the entire construction work. For the direct approach, the description of the Tutte-decomposition via total-nestedness is key. The construction from [27] can be generalised so that it can be applied to the tri-separation decomposition; we would be excited to see this happen.
@@
It would be most exciting to try to extend our work to separators of larger size, see Section 7 for an open question in this direction. Just recently, our work has been extended to separators of size 4 in [51]. Our work complements existing work on decomposing graphs along separators of arbitrary size and the corresponding theory of tangles [2, 7, 10, 11, 12, 13, 16, 22, 23, 25, 26, 29, 30, 31, 40, 41, 44, 59]. A fair share of the work on 3-connected graphs studies which substructures are ‘(in-)essential’ to 3-connectivity (in the sense that their contraction or deletion preserves 3-connectivity); this includes the work of Ando, Enomoto and Saito [3] and of Kriesell [45, 46, 47, 48, 49, 50]. The structure of 3-separations in matroids is a well-studied topic; a fundamental result in this area is the decomposition result of Oxley, Semple and Whittle [55]. It would be most natural to extend our ideas to matroids and this problem is discussed in Section 7. Also see [5]. Hopes for a generalisation to directed graphs are fuelled by recent work of Bowler, Gut, Hatzel, Kawarabayashi, Muzi and Reich [4]. Our work is related to recent work of Esperet, Giocanti and Legrand-Duchesne [32], who employ Grohe’s techniques for 3-connected graphs to prove a general decomposition result of infinite graphs without an infinite clique-minor, and our result might provide an alternative perspective. There is also a connection to twin-width, see for example the recent work of Heinrich and Raßmann [43].
@@
An overview of the proof of Theorem 1 is given in Section 0. The remainder of the paper consists of four chapters. Each chapter will feature its own comprehensive overview. In the first chapter, we introduce and prove the (Angry Tri-Separation Theorem). (1.1); this classifies the 3-connected graphs in which all nontrivial tri-separations are crossed (hence the name), and it will be the key ingredient of the proof of Theorem 1 as it deals with the special case $N(G)=\emptyset$. Applications can already be derived from the (Angry Tri-Separation Theorem)., so the first chapter also includes the proof of Applications. In the second chapter, we prove Theorem 1. The third chapter provides an outlook. Finally, the fourth chapter offers a proof of the 2-separator theorem, but phrased in the language of this paper and with a structural strengthening added, for the sake of convenience and completeness.
0 Overview of the proof
Let $G$ be a 3-connected graph, and let $N$ denote the set of totally-nested nontrivial tri-separations of $G$. By adapting standard arguments, we will find a unique mixed-tree-decomposition ${{\mathcal{T}}}=(T,(V_{t})_{t\in V(T)})$ of $G$ such that the edges of $T$ induce precisely the tri-separations in $N$. Let $\tau$ be an arbitrary torso of $N$, which can also be viewed as a torso of bag $V_{t}$ of the mixed-tree-decomposition ${{\mathcal{T}}}$. It is routine to verify that $\tau$ is 3-connected or a $K_{3}$, and that $\tau$ is a minor of $G$. So it remains to show that $\tau$ either is quasi 4-connected, a wheel, a thickened $K_{3,m}$ or $G=K_{3,m}$ with $m\geqslant 0$.
@@ 0 Overview of the proof
For the proof of (3), suppose that $\tau$ is lightly interlaced by a tri-separation $(A,B)$. Then we first note that the separator $S$ of $(A,B)$ consists of three vertices, and that $G\setminus S$ has at least four components. The four components ensure that $S$ is ‘4-connected’ in $G$, as every two vertices in $S$ can be linked by four internally vertex-disjoint paths in $G$ through the four components. The 4-connectivity of $S$ can then be used to show that every component $C$ of $G\setminus S$ determines a totally-nested tri-separation of $G$, with $C$ on one side, whose separator consists of vertices of $S$ or edges that join $C$ to $S$. The nontrivial ones amongst the totally-nested tri-separations determined by the components of $G\setminus S$ are precisely the ones that bound the torso $\tau$ (i.e., the ones that are induced by the edges of $T$ at $t$). If $G[S]$ is edgeless and all components of $G\setminus S$ have size one, then $G=K_{3,m}$ where $m$ is equal to the number of components of $G\setminus S$. Otherwise, a brief analysis shows that $\tau$ is a thickened $K_{3,m}$, where $m$ is equal to the number of components $C$ of $G\setminus S$ such that $|C|=1$ and $G[S]$ is edgeless.
Figure 1: The graph $G$ (on the right) can be constructed from a cycle $O$ (on the left) by replacing its bold edges by 2-connected graphs and adding the vertex $v$ together with its incident edges. Here $S(A,B)$ and $S(C,D)$ denote the separators of $(A,B)$ and $(C,D)$, respectively. They intersect in the vertex $v$. The cycle $O$ is highlighted in green.
@@
Figure 2: The graph $G$ together with the bold edges of $O$. For each bold edge of $O$ we indicate in grey its corresponding 2-connected replacement graph as given in the construction of $G$ from $O$. In this figure we refer to them as bags. The vertex $a$ is an endvertex of the bold edge $e$. Here the cycle $O$ has the pattern btx with regard to $e$ and $a$: the first letter b indicates that the edge $f_{1}$ on $O$ incident with $a$ aside from $e$ is bold, the second letter t indicates that the edge $f_{2}$ after that on $O$ is not bold (timid), and the letter x indicates that the edge $f_{3}$ after that can be arbitrary, bold or not. In our example, it is bold. For this pattern, there are three internally vertex-disjoint paths from $a$ to $v$: the first path $P_{1}$ connects to $v$ from the bag at $e$, the second path $P_{2}$ connects to $v$ from the bag at $f_{1}$, and the final path $P_{3}$ uses a path disjoint from $P_{2}-a$ through the bag at $f_{1}$, then traverses $f_{2}$ and eventually connects to $v$ from the bag at $f_{3}$.
@@
Having shown that the pseudo-reductions at the bold edges of $O$ are totally-nested, we know that they bound a torso $\tau^{\prime}$; this is not necessarily a torso of $N$, but of the set of pseudo-reductions. That is to say that $\tau^{\prime}$ is the torso of a mixed-tree-decomposition which is coarser than ${{\mathcal{T}}}$, so $\tau^{\prime}$ includes $\tau$, and we will later show that $\tau^{\prime}=\tau$. Using our knowledge of the structure of $G-v$ provided by $O$, and 3-connectedness of $G$, it is straightforward to show that $\tau^{\prime}$ is a wheel. So all that remains to show is that $\tau^{\prime}$ is equal to $\tau$. For this, we have to show that the pseudo-reductions at the bold edges of $O$ are precisely the tri-separations induced by the edges of $T$ at $t$. Equivalently, we have to show that no totally-nested nontrivial tri-separation of $G$ interlaces $\tau^{\prime}$. So we assume for a contradiction that some totally-nested nontrivial tri-separation $(U,W)$ of $G$ interlaces $\tau^{\prime}$. Roughly speaking, $(U,W)$ induces a mixed 2-separation of $O$. This induced mixed 2-separation cuts $O$ into two intervals. We carefully select a vertex or non-bold edge from each interval, and then add $v$ to obtain the separator of a mixed 3-separation $(E,F)$ of $G$. Then $(E,F)$ crosses $(U,W)$ by standard arguments, and with a bit of extra work we turn $(E,F)$ into a tri-separation that still crosses $(U,W)$, yielding a contradiction. Hence $\tau^{\prime}$ is equal to $\tau$. As we have already shown that $\tau^{\prime}$ is a wheel, the overview of the proof of (2) is complete.
Chapter 1
An angry theorem for tri-separations
1 Overview of this chapter
@@
This chapter provides the key ingredient for the proof of Theorem 1, which we call the (Angry Tri-Separation Theorem).. Essentially, it states that Theorem 1 holds in the special case where $N(G)$, the set of all totally-nested non-trivial tri-separations of $G$, is empty; that is: if every nontrivial tri-separation of $G$ is crossed (hence the name ‘angry’), then $G$ is either quasi 4-connected, a wheel or a $K_{3,m}$ for some $m\geqslant 3$. The (Angry Tri-Separation Theorem). extends to 3-connectivity the (Angry 2-Separation Theorem). (3.1) of Tutte, which characterises the 2-connected graphs in which every every 2-separation is crossed by some other 2-separation: these are precisely the 3-connected graphs and the cycles. We recommend having an early glance at Section 1 to see how the (Angry 2-Separation Theorem). fits into the proof of Tutte’s decomposition for 2-connectivity.
1 Statement of the (Angry Tri-Separation Theorem).
@@
Here we give all the necessary definitions to then state the (Angry Tri-Separation Theorem). (1.1). An (oriented) mixed-separation of a graph $G$ is an ordered pair $(A,B)$ such that $A\cup B=V(G)$ and both $A\setminus B$ and $B\setminus A$ are non-empty. We call $A$ and $B$ the sides of $(A,B)$. The separator of $(A,B)$ is the disjoint union of the vertex set $A\cap B$ and the edge set $E(A\setminus B,B\setminus A)$. We denote the separator of $(A,B)$ by $S(A,B)$. The order of $(A,B)$ is the size $|S(A,B)|$ of its separator. A mixed-separation of order $k$ for $k\in{\mathbb{N}}$ is called a mixed $k$-separation for short. The separator of a mixed $k$-separation is a mixed $k$-separator. A mixed-separation $(A,B)$ of $G$ with no edges in its separator is called a separation of $G$. Separations of order $k$ are called $k$-separations and their separators are called $k$-separators.
@@
Figure 3: Separators of nontrivial tri-separations of the 4-wheel.
@@
Every nontrivial tri-separation in the 4-wheel, as found in Section 1, is crossed by another nontrivial tri-separation; see Figure 3. By contrast, every 3-cut with both sides of size at least two in a 3-connected graph determines a totally-nested nontrivial tri-separation (we will see this in Section 4).
@@
Theorem 1.1 (Angry Tri-Separation Theorem).
@@
2 Organisation of this chapter
In Section 2, we explore nested sets of tri-separations from a mixed-tree-decomposition perspective. In Section 3, we collect useful properties of tri-separations. In Section 4, we introduce tools that allow us to systematically study how tri-separations can cross. In Section 5, we deduce Applications from the (Angry Tri-Separation Theorem).. In Section 6, we employ the tools from the previous section to find necessary conditions for when a tri-separation is totally nested. In Section 7, we recall the 2-separation-version of Theorem 1, as we will need it in the proof of the (Angry Tri-Separation Theorem).. In Section 8, we will see why the 2-separation-version of Theorem 1 is helpful for finding totally-nested nontrivial tri-separations. In Section 9 and Section 10, we put together the tools developed in the previous sections and we complete the proof of the (Angry Tri-Separation Theorem)., where the former deals with special cases and the latter solves the general case.
2 Mixed-tree-decompositions
  1. (M1)
    $V(G)=\bigcup_{t\in T}V_{t}$;
  2. (M2)
We refer to the vertex sets $V_{t}$ as bags.
[…]
3 Properties of tri-separations